a rectangle has a perimeter of 56 inches and a base of 10 inches what is the height

Answer by JulietG(1812)

2L + 2W = perimeter (think of a perimeter as a fence around the rectangle)
W = 20
2L + 2(10) = 56
2L + 20 = 56
subtract 20 from each side
2L = 36
Divide each side by 2
L = 18